OpenStudy (anonymous):

volume bounded by xy plane,the cylinder (x^2+y^2=1)and the plane (x+y+z=1)

OpenStudy (zarkon):

\[\int_0^{2\pi} r (-r \sin (\theta )-r \cos (\theta )+1)d\theta=2\pi r\]

OpenStudy (anonymous):

\[x = r \cos(\theta) \\ y = r \sin(\theta)\\ dV= r dr d\theta\\ \int_o^1\int_0^{2\pi} r (-r \sin (\theta )-r \cos (\theta )+1)d\theta dr=\\ \int_o^1 2 r \pi dr=\pi \]